# Practicalities of State-Dependent and Threshold Delay Differential Equations

### v1

@@ Acknowledgments
$u_{t}(\theta)=u(t+\theta),\quad\theta\in[-\tau,0],$ (1.5)
@@
Figure 1.1: Illustration of (1.6) the initial condition for an RFDE, and a function segment $u_{t}$ for $t=t_{0}+\tau/2$ illustrating that $u_{t}$ is not necessarily continuously differentiable, even when $\varphi$ is smooth.
@@
So far, we have only considered constant delays, but in many applications the delay can depend on the state of the system; $\tau(u(t))$, or in an RFDE $\tau(u_{t})$. The analysis of RFDEs with state-dependent delays becomes very complicated because $C$ being a space of only continuous functions then results in $F$ not being a Lipschitz function, and standard results for smooth functions cannot be applied directly. See the review Hartung et al. (2006) for a description of the theoretical issues and approaches for their resolution.
@@ 1 DDE Model Examples
Figure 1.2: Schematic representation of the classical $G_{0}$ model for HSCs. The proliferating phase of the cell cycle is divided between 4 subphases: gap one $G_{1}$, synthesis $S$, gap two $G_{2}$, and mitosis $M$. Cells in the resting phase (gap zero $G_{0}$), may differentiate with rate $\kappa$ or entry the cell cycle with rate $\beta(Q)$. Cells in the proliferation phase may be lost by apoptosis with rate $\gamma$, otherwise they re-enter the resting phase after mitosis, $\tau$ time units after they left the resting phase. (© Society for Industrial and Applied Mathematics (SIAM); Reproduced from De Souza and Humphries (2019) with permission)
@@
The general DDE (1.7) will be a state-dependent delay DDE, if $\tau$ instead of being a constant delay, is allowed to depend on the state. The simplest way to do this is to make $\tau$ a function of the state-variable at time $t$, so $\tau=\tau(u(t))>0$. We will the explore the dynamics of the DDE (1.22) with two such state-dependent delays in Section 2.
@@
for a given non-negative function $V$ and positive constant $a$, will also be a state-dependent delay DDE. We remark that the DDE (1.7) with delay $\tau$ defined by the threshold condition (1.21) is no longer a discrete delay DDE, as the evolution of the solution depends on the delay $\tau$ which itself depends integral of $V(u(s))$ for $s\in[t-\tau,t]$. Thus (1.7),(1.21) is properly regarded as an RFDE, but one where $\dot{u}(t)$ only depends on the values of $u(t)$ and $u(t-\tau)$, but where $\tau$ itself is defined the function $u_{t}$. This kind of halfway house between discrete delays and RFDEs will turn out to make threshold problems more tractable than general RFDEs.
@@ 2 Why model with state-dependent delays?
It has two state-dependent delays $\tau_{j}(u(t))=a_{i}+c_{j}u(t)$ which become constant delays when $c_{1}=c_{2}=0$, and are linearly state-dependent otherwise.
@@
Figure 1.4: Solutions of the IVP (1.23) with constant initial functions $\varphi(t)=u_{0}$ for several different values of $u_{0}$ with other parameters $1=a_{1}=\gamma<\kappa_{2}=\kappa_{1}=a_{2}=2$, $c_{1}=0.5$, $c_{2}=0.4$. The dashed line at u=2u=-2 indicates where $u(t)=-a_{1}/c_{1}$ and $\tau_{1}=0$. (© American Institute of Mathematical Sciences (AIMS); Reproduced from Humphries et al. (2012) with permission)
@@
Figure 1.5: Solutions of (1.22) projected onto $(u(t),u(t-a_{1}),u(t-a_{2}))$-space. Parameter values are $\gamma=4.75>\kappa_{2}=3.0$, $a_{1}=1.3$, $a_{2}=6.0$ and $c_{1}=c_{2}=1.0$. (a) With $\kappa_{1}=4.44$ a quasi-periodic torus is observed. (b) For $\kappa_{1}=6.93$, a 1:4 phase-locked torus-like object (grey), with embedded stable (blue) and saddle (red) periodic orbits. (©SIAM; Reproduced from Calleja et al. (2017) with permission)
@@ 3 A scalar DDE with threshold delay
Then since $0<V_{\textit{min}}\leqslant V(u(t))\leqslant V_{\textit{max}}$ it follows that process described by (2.1) always proceeds to completion without stalling. That is, the threshold delay $\tau=\tau(u_{t})$ is also bounded and bounded away from zero with
@@ 5 The Velocity Ratio In Threshold Delay Models
Now, suppose the conveyor belt is very long, and has adjustable speed. The bags are still placed on the belt at a rate of one a second, but the velocity of the conveyor belt is doubled before the first bag arrives. Now, once the bags start to arrive, they will do so at a rate of the two every second, until such time as all the bags that were placed on the belt when it was moving slowly have arrived.
@@
Figure 2.2: (a) Illustrating the characteristics of (2.14) with constant velocity $V$ and the relationship (2.21) between the initial function $\varphi$ for (2.2) and the boundary condition $\psi$ for (2.14). (b) Illustration of maturation at with non-constant ageing velocity.
@@
The state-dependent delays also need to be well behaved to apply the theorems, in particular we require $\tau\geqslant 0$. We remark that for specific examples, such as the linear state-dependent delays considered in Section 2 this can be ensured either by imposing parameter constraints, or by modifying the definition of the delays suitability. The threshold delay (2.1) can never become advanced, and the differentiated form (2.7) allows the theory to be applied. It also follows from (2.6) that $t-\tau$ is a monotonically increasing function of $t$ along any solution trajectory.
@@
Another issue with an infinite dimensional phase space, is how to represent solutions visually. For ODEs in two or three space dimensions its usual to display solutions as a parametrized curves in phase space, but for a DDE we will need to project the phase portrait into finite dimensions to display it. Its not clear how to best do this. When $u(t)$ is analytic, the solution will be equal to its Taylor series, so the most obvious and natural projection would be to project the function $u(t)$ onto the coefficients of its Taylor series evaluated at $t$, and so plot the first few coordinates of $(u(t),\dot{u}(t),\ddot{u}(t),\ldots)$ or $(u(t),\dot{u}(t)/1!,\ddot{u}(t)/2!,\ldots)$ as parametrized curves. But this is problematical because the breaking points mean that generally solutions are not analytic, and even when they are the derivatives may not scale nicely.
@@ 6 Invariant Sets and Bifurcations
If $A$ is an invariant set, then there is a complete orbit passing through every point of $A$; so for $u\in A$, we can define $S(t)u\in A$ for all $t\in\mathbb{R}$. This then allows for the definition of $\alpha$- and $\omega$-limit sets and attractors. To show that $\omega$-limit sets are compact and non-empty is not as simple as for the finite-dimensional case of ODEs, but for can be done using the Arzelà-Ascoli Theorem; see Chapter 5 of Smith (2011), where this is done for RFDEs.
@@
We will be interested in the stability of steady states and other invariant sets, and the bifurcations that occur as parameters are varied. We will obtain the stability of steady states by linearizing around then. In contrast to partial differential equations for which there is no general principle of linearized stability, DDEs and RFDEs, although defining infinite-dimensional dynamical systems are better behaved, and do have possess a principle of linearized stability Smith (2011). Consequently a steady state will be stable if all the characteristic values $\lambda$ have negative real part, and unstable if one or more characteristic values $\lambda$ have positive real part. Floquet theory generalises this technique to periodic orbits.
On a periodic orbit of (3.1), or indeed on any complete orbit contained in a compact invariant set, the smoothing property described in Section 3 ensures that the solution is infinitely differentiable when $f$ is infinitely differentiable. When the delay(s) are constant, it then follows from the results of Nussbaum (1973) that the solution is analytic when $f$ is analytic. However, Mallet-Paret and Nussbaum (2014) showed that if discrete delays are allowed to be time or state dependent then solutions may not be analytic for all $t$. Krisztin (2014) on the other hand showed that with a threshold delay defined by (2.1) the bounded solutions of (3.1) are analytic when $f$ and $V$ are both analytic.
@@
where $L$ is the Frechet derivative of $F$ evaluated at the constant function $u_{t}=u^{*}\in\mathbb{R}^{d}$. If you do not know what this is do not worry; for a state-dependent delay, $F$ is not Lipschitz and so not differentiable and this does not work anyway. It was only in the 21st century that a rigorous theory was developed for the linearisation of state-dependent delay problems. However, it involves working in a subspace of $C$ on which all elements $\varphi$ are $C^{1}$ and satisfy $\varphi^{\prime}(0)=F(\varphi)$, and essentially differentiating there, before extending the result to $C$ Walther (2003); Hartung et al. (2006). In the hope of reaching readers for whom such analysis arguments are opaque, we will show how to linearise heuristically for both discrete state-dependent delays and threshold delays, and also illustrate the pit holes to avoid falling down.
@@ 7 Linearisation for Discrete Delays
using the Taylor series in two variables for $f(u,v)$. Dropping the higher order terms (arising from higher derivatives of $f$) we obtain the linearized equation
@@
for constants $\mu\in\mathbb{R}$, $\sigma\in\mathbb{R}$. Positing $u(t)=e^{\lambda t}$ gives transcendental characteristic equation, $\Delta(\lambda)=0$, where $\Delta$ is the characteristic function
@@
DDEs posed for $u(t)\in\mathbb{R}^{d}$ can be linearized similarly. Consider
@@
where $A_{0}=f_{u}$ and $A_{j}=f_{v_{j}}$ are $d\times d$ matrices evaluated at the steady-state (essentially a Jacobian matrix for each delay). There is nontrivial solution $u(t)=e^{\lambda t}\underline{v}\in\mathbb{R}^{d}$ with $\Delta(\lambda)\underline{v}=0$ if
@@
Linearization with State-Dependent Delays
@@
$\dot{u}(t)=f(u(t),u(t-\tau(u_{t}))),$ (3.15)
The original heuristic method to linearize about a steady state $u^{*}$ is to freeze the delay $\tau(u_{t})$ at its steady state value $\tau(u^{*})$ and then linearize the resulting constant delay DDE
@@
$\dot{u}(t)=-\gamma u(t)-\sum_{i=1}^{2}\kappa_{i}u(t-a_{i})$ (3.17)
Since (3.17) is already linear in this case, this completes the linearization.
An analytic approach using the semiflow in an appropriate Banach space developed in Walther (2003); Hartung et al. (2006) (see also Cooke and Huang (1996); Gyori and Hartung (2007)) leads to the same linearisation when the delay $\tau(u_{t})$ is a discrete delay, for example when $\tau(u_{t}))=\tau(u_{t}(0)))=\tau(u(t))$, but shows that freezing the delay leads to an incorrect linearisation for distributed delays, such as threshold delays. Importantly, the work of Walther (2003); Hartung et al. (2006) establishes the principal of linearized stability for DDEs with state-dependent delays.
@@
8 Linearization of Threshold Delays
@@
Linearization by Perturbation from Linear Dynamics
@@
Linearization by Expansion
An alternative derivation of (3.30), is to first linearize the system (2.2),(2.1), and then determine the dynamics of the linearized system. Begin by defining new variables $\tilde{u}(t)$ and $\tilde{\tau}(t)$ by
@@
Now combining (8) and (3.32) and neglecting the higher order terms, we get the linearized system
@@
Linearization with differentiation of the threshold condition
In Section 3 we proposed to replace the threshold condition (2.1) by its differentiated form (2.7) to convert the threshold delay system (2.2),(2.1) into the discrete delay system (2.1),(2.7). To linearize (2.7) around a steady state, substitute (3.31) into (2.7) to obtain
@@ 9 IVPs with Discrete State-Dependent Delays
The interpolant $u^{h^{\!}}(t_{n}+\theta h_{n})$ in (4.4) allows us to extent $u^{h^{\!}}$ in (4.3) to the interval $t\in[t_{0},t_{n+1}]$, while equation (4.5) defines explicit interpolants which are only used in the computation of the $i$-th stage value $k_{n,i}$ if there is overlapping (if $t-\tau(t_{n}+c_{i}h_{n},U_{n,i})>t_{n}$).
@@
for integer $N$ and constant $\lambda\in[0,1)$ so that the breaking point always falls at the same fraction $\lambda$ in the $(N+1)$-st step, for any choice of the integer $N$. The error between the the numerical and exact solutions as functions is measured using the $L^{\infty}$-norm
@@
Several issues are at play in the example of Figure 4.1(d). Firstly, because breaking point locations depend on the solution, the breaking points $\xi_{j}^{h}$ of the numerical solution $u^{h^{\!}}(t)$ and the breaking points $\xi_{j}$ of the exact solution $u(t)$ will not correspond exactly, and it is $\xi_{j}^{h}$, and not $\xi_{j}$ that should be included in the computational mesh. Secondly even if $\xi_{j}^{h}-\xi_{j}=\mathcal{O}(h^{p})$, for an order 0 breaking point $\xi^{h}_{j}$, there will be an $\mathcal{O}(1)$ difference between $f(t,u^{h^{\!}}(t),t-\tau(t,u^{h^{\!}}(t)))$ for $t$ values close to $\xi^{h}_{j}$ such that $t-\tau(t,u^{h^{\!}}(t))$ falls either side of the previous order 1-1 breaking point. In this case an arbitrary small error in determining the breaking point can generate a $\mathcal{O}(1)$ error in an evaluation of $f$. Thus, it is not sufficient just to detect breaking points to accuracy (4.9), but how these breaking points are handled in the evaluation of the numerical solution with be crucial for maintaining the order of the method. In particular, it will be necessary to ensure that function values are never evaluated from the wrong side of a previous breaking point.
@@
$\left.\begin{aligned} \dot{u}(t)&=f(t,u(t),u(t-\tau)),&&t\in[t_{n},t_{n+1}],\\ u(t_{n})&=u^{h^{\!}}(t_{n}),\\ u(t-\tau)&=u^{h}_{-\!}(t-\tau),&&t-\tau\leqslant\xi_{i}^{h},\\ u(t-\tau)&=u^{h^{\!}}_{+\!}(t-\tau),&&t-\tau\geqslant\xi_{i}^{h},\end{aligned}\right\}$ (4.17)
@@
shown in Figure 4.3. In (4.18) we analytically continue $u^{h}_{-\!}(t-\tau)$ across the breaking point $\xi_{i}^{h}$ to to create a modified problem without a breaking point. The new problem (4.18) is smooth over the step $t\in[t_{n},t_{n+1}]$ and has an identical solution to (4.17) for $t\in[t_{n},\xi_{j}^{h}]$. Thus we solve (4.18) for $t\in[t_{n},t_{n+1}]$, and use this solution to approximate the new breaking point $\xi_{j}^{h}$ to the full accuracy of the FCRK method. Once $\xi_{j}^{h}$ is found, the solution of (4.18) for $t\in[t_{n},\xi_{j}]$ can be used to define $u^{h^{\!}}(t)$ for $t\in[t_{n},\xi_{j}^{h}],$ and we proceed to the next step defining $t_{n+1}:=\xi_{j}^{h}$.
@@
Figure 4.4(c) and (d) show the 4 FCRK methods applied to (4.12) over the time interval $t\in[0,1]$, for which there is a breaking point of order 0 at $\xi_{j}\in(0,1)$. No breaking point detection is applied for FCRK1, while for FCRK2 and FCRK3 the breaking point is approximated by $\xi_{j}^{p}$ defined by (4.27). For FCRK4 the Secant approximation $\xi_{j}^{s}$ defined by (4.30) is used to approximate $\xi_{j}$. Recall from Figure 4.1(d) that without breaking point detection all of the methods displayed only first order convergence, even when the exact breaking point $\xi_{j}$ was included in the computational mesh. Figure 4.4(c) and (d) shows that applying our breaking point detection algorithm, all of the methods retain their full order, and also find the breaking point to the full order of accuracy of the method.

### v2

@@ Acknowledgments
$u_{t}(\theta)=u(t+\theta),\quad\theta\in[-\tau,0].$ (1.5)
@@
Figure 1.1: Illustration of (1.6), the initial condition for an RFDE, and a function segment $u_{t}$ for $t=t_{0}+\tau/2$ illustrating that $u_{t}$ is not necessarily continuously differentiable, even when $\varphi$ is smooth.
@@
So far, we have only considered constant delays, but in many applications the delay can depend on the state of the system, $\tau(u(t))$, or in an RFDE $\tau(u_{t})$. The analysis of RFDEs with state-dependent delays becomes very complicated because $C$ being a space of only continuous functions then results in $F$ not being a Lipschitz function, and standard results for smooth functions cannot be applied directly. See the review Hartung et al. (2006) for a description of the theoretical issues and approaches for their resolution.
@@ 1 DDE Model Examples
Figure 1.2: Schematic representation of the classical $G_{0}$ model for HSCs. The proliferating phase of the cell cycle is divided between 4 subphases: gap one $G_{1}$, synthesis $S$, gap two $G_{2}$, and mitosis $M$. Cells in the resting phase (gap zero $G_{0}$), may differentiate with rate $\kappa$ or entry the cell cycle with rate $\beta(Q)$. Cells in the proliferation phase may be lost by apoptosis with rate $\gamma$, otherwise they re-enter the resting phase after mitosis, $\tau$ time units after they left the resting phase. (© Society for Industrial and Applied Mathematics (SIAM); Reproduced from De Souza and Humphries (2019) with permission.)
@@
The general DDE (1.7) will be a state-dependent delay DDE, if $\tau$ instead of being a constant delay, is allowed to depend on the state. The simplest way to do this is to make $\tau$ a function of the state-variable at time $t$, so $\tau=\tau(u(t))>0$. In Section 2, we will explore the dynamics of the DDE (1.22) with two state-dependent delays of this form.
@@
for a given non-negative function $V$ and positive constant $a$, will also be a state-dependent delay DDE. We remark that the DDE (1.7) with delay $\tau$ defined by the threshold condition (1.21) is no longer a discrete delay DDE, as the evolution of the solution depends on the delay $\tau$ which itself depends integral of $V(u(s))$ for $s\in[t-\tau,t]$. Thus (1.7),(1.21) is properly regarded as an RFDE, but one where $\dot{u}(t)$ only depends on the values of $u(t)$ and $u(t-\tau)$, but where $\tau$ itself is defined as a function of $u_{t}$. This kind of halfway house between discrete delays and RFDEs will turn out to make threshold problems more tractable than general RFDEs.
@@ 2 Why model with state-dependent delays?
It has two state-dependent delays $\tau_{j}(u(t))=a_{j}+c_{j}u(t)$ which become constant delays when $c_{1}=c_{2}=0$, and are linearly state-dependent otherwise.
@@
Figure 1.4: Solutions of the IVP for (1.23) with constant initial functions $\varphi(t)=u_{0}$ for several different values of $u_{0}$ with other parameters $1=a_{1}=\gamma<\kappa_{2}=\kappa_{1}=a_{2}=2$, $c_{1}=0.5$, $c_{2}=0.4$. The dashed line at u=2u=-2 indicates where $u(t)=-a_{1}/c_{1}$ and $\tau_{1}=0$. (© American Institute of Mathematical Sciences (AIMS); Reproduced from Humphries et al. (2012) with permission.)
@@
Figure 1.5: Solutions of (1.22) projected onto $(u(t),u(t-a_{1}),u(t-a_{2}))$-space. Parameter values are $\gamma=4.75>\kappa_{2}=3.0$, $a_{1}=1.3$, $a_{2}=6.0$ and $c_{1}=c_{2}=1.0$. (a) With $\kappa_{1}=4.44$ a quasi-periodic torus is observed. (b) For $\kappa_{1}=6.93$, a 1:4 phase-locked torus-like object (grey), with embedded stable (blue) and saddle (red) periodic orbits. (©SIAM; Reproduced from Calleja et al. (2017) with permission.)
@@ 3 A scalar DDE with threshold delay
Then since $0<V_{\textit{min}}\leqslant V(u(t))\leqslant V_{\textit{max}}$ it follows that the process described by (2.1) always proceeds to completion without stalling. That is, the threshold delay $\tau=\tau(u_{t})$ is also bounded and bounded away from zero with
@@ 5 The Velocity Ratio In Threshold Delay Models
Now, suppose the conveyor belt is very long, and has adjustable speed. The bags are still placed on the belt at a rate of one a second, but the velocity of the conveyor belt is doubled before the first bag arrives. Once the bags start to arrive, they will do so at a rate of two every second, until such time as all the bags that were placed on the belt when it was moving slowly have arrived.
@@
Figure 2.2: (a) Illustrating the characteristics of (2.14) with constant velocity $V$ and the relationship (2.21) between the initial function $\varphi$ for (2.2) and the boundary condition $\psi$ for (2.14). (b) Illustration of maturation with non-constant ageing velocity.
@@
The state-dependent delays also need to be well behaved to apply the theorems, in particular we require $\tau\geqslant 0$. We remark that for specific examples, such as the linear state-dependent delays considered in Section 2 this can be ensured either by imposing parameter constraints, or by modifying the definition of the delays suitably. The threshold delay (2.1) can never become advanced, and the differentiated form (2.7) allows the theory to be applied. It also follows from (2.6) that $t-\tau$ is a monotonically increasing function of $t$ along any solution trajectory.
@@
Another issue with an infinite dimensional phase space, is how to represent solutions visually. For ODEs in two or three space dimensions its usual to display solutions as parametrised curves in phase space, but for a DDE we will need to project the phase portrait into finite dimensions to display it. Its not clear how to best do this. When $u(t)$ is analytic, the solution will be equal to its Taylor series, so the most obvious and natural projection would be to project the function $u(t)$ onto the coefficients of its Taylor series evaluated at $t$, and so plot the first few coordinates of $(u(t),\dot{u}(t),\ddot{u}(t),\ldots)$ or $(u(t),\dot{u}(t)/1!,\ddot{u}(t)/2!,\ldots)$ as parametrised curves. But this is problematical because the breaking points mean that generally solutions are not analytic, and even when they are the derivatives may not scale nicely.
@@ 6 Invariant Sets and Bifurcations
If $A$ is an invariant set, then there is a complete orbit passing through every point of $A$; so for $u\in A$, we can define $S(t)u\in A$ for all $t\in\mathbb{R}$. This then allows for the definition of $\alpha$- and $\omega$-limit sets and attractors. To show that $\omega$-limit sets are compact and non-empty is not as simple as for the finite-dimensional case of ODEs, but can be done using the Arzelà-Ascoli Theorem; see Chapter 5 of Smith (2011), where this is done for RFDEs.
@@
We will be interested in the stability of steady states and other invariant sets, and the bifurcations that occur as parameters are varied. We will obtain the stability of steady states by linearizing around then. In contrast to partial differential equations for which there is no general principle of linearised stability, DDEs and RFDEs, although defining infinite-dimensional dynamical systems are better behaved, and do have possess a principle of linearised stability Smith (2011). Consequently a steady state will be stable if all the characteristic values $\lambda$ have negative real part, and unstable if one or more characteristic values $\lambda$ have positive real part. Floquét theory generalises this technique to periodic orbits.
On a periodic orbit of (3.1), or indeed on any complete orbit contained in a compact invariant set, the smoothing property described in Section 3 ensures that the solution is infinitely differentiable when $f$ is infinitely differentiable. When the delay is constant, it then follows from the results of Nussbaum (1973) that the solution is analytic when $f$ is analytic. However, Mallet-Paret and Nussbaum (2014) showed that if discrete delays are allowed to be time or state dependent then solutions may not be analytic for all $t$. Krisztin (2014) on the other hand showed that with a threshold delay defined by (2.1) the bounded solutions of (3.1) are analytic when $f$ and $V$ are both analytic.
@@
where $L$ is the Fréchet derivative of $F$ evaluated at the constant function $u_{t}=u^{*}\in\mathbb{R}^{d}$. If you do not know what this is do not worry; for a state-dependent delay, $F$ is not Lipschitz and so not differentiable and this does not work anyway. It was only in the 21st century that a rigorous theory was developed for the linearisation of state-dependent delay problems. However, it involves working in a subspace of $C$ on which all elements $\varphi$ are $C^{1}$ and satisfy $\varphi^{\prime}(0)=F(\varphi)$, and essentially differentiating there, before extending the result to $C$ Walther (2003); Hartung et al. (2006). In the hope of reaching readers for whom such analysis arguments are opaque, we will show how to linearise heuristically for both discrete state-dependent delays and threshold delays, and also illustrate the pit holes to avoid falling down.
@@ 7 Linearisation for Discrete Delays
using the Taylor series in two variables for $f(u,v)$. Dropping the higher order terms (arising from higher derivatives of $f$) we obtain the linearised equation
@@
for constants $\mu\in\mathbb{R}$, $\sigma\in\mathbb{R}$. Positing $u(t)=e^{\lambda t}$ gives a transcendental characteristic equation, $\Delta(\lambda)=0$, where $\Delta$ is the characteristic function
@@
DDEs posed for $u(t)\in\mathbb{R}^{d}$ can be linearised similarly. Consider
@@
where $A_{0}=f_{u}$ and $A_{j}=f_{v_{j}}$ are $d\times d$ matrices evaluated at the steady-state (essentially a Jacobian matrix for each delay). There is a nontrivial solution $u(t)=e^{\lambda t}\underline{v}\in\mathbb{R}^{d}$ with $\Delta(\lambda)\underline{v}=0$ if
@@
Linearisation with State-Dependent Delays
@@
$\dot{u}(t)=f(u(t),u(t-\tau(u_{t}))).$ (3.15)
The original heuristic method to linearise about a steady state $u^{*}$ is to freeze the delay $\tau(u_{t})$ at its steady state value $\tau(u^{*})$ and then linearise the resulting constant delay DDE
@@
$\dot{u}(t)=-\gamma u(t)-\sum_{i=1}^{2}\kappa_{i}u(t-a_{i}).$ (3.17)
Since (3.17) is already linear in this case, this completes the linearisation.
An analytic approach using the semiflow in an appropriate Banach space developed in Walther (2003); Hartung et al. (2006) (see also Cooke and Huang (1996); Gyori and Hartung (2007)) leads to the same linearisation when the delay $\tau(u_{t})$ is a discrete delay, for example when $\tau(u_{t}))=\tau(u_{t}(0)))=\tau(u(t))$, but shows that freezing the delay leads to an incorrect linearisation for distributed delays, such as threshold delays. Importantly, the work of Walther (2003); Hartung et al. (2006) establishes the principle of linearised stability for DDEs with state-dependent delays.
@@
8 Linearisation of Threshold Delays
@@
Linearisation by Perturbation from Linear Dynamics
@@
Linearisation by Expansion
An alternative derivation of (3.30), is to first linearise the system (2.2),(2.1), and then determine the dynamics of the linearised system. Begin by defining new variables $\tilde{u}(t)$ and $\tilde{\tau}(t)$ by
@@
Now combining (8) and (3.32) and neglecting the higher order terms, we get the linearised system
@@
Linearisation with differentiation of the threshold condition
In Section 3 we proposed to replace the threshold condition (2.1) by its differentiated form (2.7) to convert the threshold delay system (2.2),(2.1) into the discrete delay system (2.1),(2.7). To linearise (2.7) around a steady state, substitute (3.31) into (2.7) to obtain
@@ 9 IVPs with Discrete State-Dependent Delays
The interpolant $u^{h^{\!}}(t_{n}+\theta h_{n})$ in (4.4) allows us to extend $u^{h^{\!}}$ in (4.3) to the interval $t\in[t_{0},t_{n+1}]$, while equation (4.5) defines explicit interpolants which are only used in the computation of the $i$-th stage value $k_{n,i}$ if there is overlapping (if $t-\tau(t_{n}+c_{i}h_{n},U_{n,i})>t_{n}$).
@@
for integer $N$ and constant $\lambda\in[0,1)$ so that the breaking point always falls at the same fraction $\lambda$ in the $(N+1)$-st step, for any choice of the integer $N$. The error between the numerical and exact solutions as functions is measured using the $L^{\infty}$-norm
@@
Several issues are at play in the example of Figure 4.1(d). Firstly, because breaking point locations depend on the solution, the breaking points $\xi_{j}^{h}$ of the numerical solution $u^{h^{\!}}(t)$ and the breaking points $\xi_{j}$ of the exact solution $u(t)$ will not correspond exactly, and it is $\xi_{j}^{h}$, and not $\xi_{j}$ that should be included in the computational mesh. Secondly even if $\xi_{j}^{h}-\xi_{j}=\mathcal{O}(h^{p})$, for an order 0 breaking point $\xi^{h}_{j}$, there will be an $\mathcal{O}(1)$ difference between $f(t,u^{h^{\!}}(t),t-\tau(t,u^{h^{\!}}(t)))$ for $t$ values close to $\xi^{h}_{j}$ such that $t-\tau(t,u^{h^{\!}}(t))$ falls either side of the previous order 1-1 breaking point. In this case an arbitrary small error in determining the breaking point can generate a $\mathcal{O}(1)$ error in an evaluation of $f$. Thus, it is not sufficient just to detect breaking points to accuracy (4.9), but how these breaking points are handled in the evaluation of the numerical solution will be crucial for maintaining the order of the method. In particular, it will be necessary to ensure that function values are never evaluated from the wrong side of a previous breaking point.
@@
$\left.\begin{aligned} \dot{u}(t)&=f(t,u(t),u(t-\tau)),&&t\in[t_{n},t_{n+1}],\\ u(t_{n})&=u^{h^{\!}}(t_{n}),\\ u(t-\tau)&=u^{h}_{-\!}(t-\tau),&&t-\tau\leqslant\xi_{i}^{h},\\ u(t-\tau)&=u^{h^{\!}}_{+\!}(t-\tau),&&t-\tau\geqslant\xi_{i}^{h}.\end{aligned}\right\}$ (4.17)
@@
shown in Figure 4.3. In (4.18) we analytically continue $u^{h}_{-\!}(t-\tau)$ across the breaking point $\xi_{i}^{h}$ to create a modified problem without a breaking point. The new problem (4.18) is smooth over the step $t\in[t_{n},t_{n+1}]$ and has an identical solution to (4.17) for $t\in[t_{n},\xi_{j}^{h}]$. Thus we solve (4.18) for $t\in[t_{n},t_{n+1}]$, and use this solution to approximate the new breaking point $\xi_{j}^{h}$ to the full accuracy of the FCRK method. Once $\xi_{j}^{h}$ is found, the solution of (4.18) for $t\in[t_{n},\xi_{j}]$ can be used to define $u^{h^{\!}}(t)$ for $t\in[t_{n},\xi_{j}^{h}],$ and we proceed to the next step defining $t_{n+1}:=\xi_{j}^{h}$.
@@
Figure 4.4(c) and (d) show the 4 FCRK methods applied to (4.12) over the time interval $t\in[0,1]$, for which there is a breaking point of order 0 at $\xi_{j}\in(0,1)$. No breaking point detection is applied for FCRK1, while for FCRK2 and FCRK3 the breaking point is approximated by $\xi_{j}^{p}$ defined by (4.27). For FCRK4 the Secant approximation $\xi_{j}^{s}$ defined by (4.30) is used to approximate $\xi_{j}$. Recall from Figure 4.1(d) that without breaking point detection all of the methods displayed only first order convergence, even when the exact breaking point $\xi_{j}$ was included in the computational mesh. Figures 4.4(c) and (d) show that applying our breaking point detection algorithm, all of the methods retain their full order, and also find the breaking point to the full order of accuracy of the method.